# NoRA: Breaking the Linear Ceiling of Low-Rank Adaptation via Manifold Expansion

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) dominates parameter-efficient fine-tuning (PEFT). However, it faces a critical “linear ceiling” in complex reasoning tasks: simply increasing the rank yields diminishing returns due to intrinsic linear constraints. We introduce NoRA (Non-linear Rank Adaptation), a weight-level parallel adapter that injects SiLU gating and structural dropout to induce manifold expansion. On the SlimOrca benchmark, NoRA breaks this linear barrier: NoRA remarkably at rank 64 (PPL 3.89) outperforms LoRA at rank 512 (PPL 3.90), demonstrating superior spectral efficiency. This advantage generalizes to mathematical reasoning, where NoRA achieves a perplexity of 1.97 on MathInstruct, significantly surpassing LoRA’s saturation point of 2.07. Mechanism analysis via Singular Value Decomposition (SVD) confirms that NoRA activates the dormant tail of the singular value spectrum, effectively preventing the rank collapse observed in linear methods.
@@ 1 Introduction
To shatter this ceiling, we introduce NoRA (Non-linear Rank Adaptation). NoRA marks a paradigm shift from linear subspace optimization to non-linear manifold deformation. By injecting SiLU gating and structural dropout directly into a fine-grained, parallel adapter architecture, NoRA unlocks the high-dimensional expressivity required for complex reasoning.
@@
  • Architecture: We propose NoRA, a fine-grained, weight-level parallel adapter that integrates non-linear gating to capture complex functional updates beyond linear approximations.
  • Empirical Scaling: We demonstrate that NoRA breaks the linear ceiling. On the large-scale SlimOrca benchmark [6], NoRA continues to scale, with NoRA at rank 64 significantly outperforming LoRA at Rank 512.
@@
  • Theoretical Mechanism: Through Singular Value Decomposition (SVD) analysis, we provide a spectral proof of manifold expansion. We show that NoRA activates the dormant tail of the singular value spectrum, preventing the rank collapse that constrains linear methods.
2 NoRA: Non-linear Rank Adaptation
We introduce NoRA, a methodology that reconciles the efficiency of parameter-efficient fine-tuning with the high-dimensional expressivity required for complex reasoning.
@@ 2.1 Preliminaries: The Linear Confinement
2.2 The NoRA Architecture
NoRA breaks this linear constraint. We retain the parallel bottleneck structure to preserve synchronization with the main branch but inject non-linearity to unlock high-rank capacity, as illustrated in Figure 1 (bottom-right). Formally, NoRA is defined as:
@@ 2.2 The NoRA Architecture
Figure 1: Architectural evolution of adaptation methods. Top Row (Module-Level): Traditional adapters operate on the output of the entire attention block. (Top-left) Serial Adapters create latency bottlenecks. (Top-right) Parallel Adapters improve efficiency but lack fine-grained control over internal projections. Bottom Row (Weight-Level): (Bottom-left) LoRA injects linear updates ($\Delta W=BA$) directly into $W_{q}$ and $W_{v}$. (Bottom-right) NoRA (Ours) operates at the same fine-grained level but introduces non-linear modeling via SiLU gating ($\sigma$) and structural Dropout ($\mathcal{D}$). This architecture combines the efficiency of weight-level injection with the high-capacity expressivity of non-linear manifolds.
@@
As shown in Figure 1, a critical distinction between NoRA and parallel adapters lies in the insertion point. Unlike traditional Parallel Adapters (top-right), which operate at the module level by processing the aggregate output of an attention block, NoRA operates at the weight level (bottom-right). By injecting updates directly into the internal query ($W_{q}$) and value ($W_{v}$) projections, NoRA fundamentally alters the internal feature dynamics of the attention mechanism rather than simply correcting its output.
@@
Linearity forces the adapter to process all input features uniformly. By introducing SiLU ($\sigma(x)=x\cdot\text{sigmoid}(x)$), NoRA gains a gating mechanism. This non-linearity allows the adapter to selectively suppress noise or amplify specific feature directions in the latent space, approximating complex decision boundaries that linear low-rank updates cannot represent.
@@
In standard PEFT, dropout is often omitted. In NoRA, we utilize dropout $\mathcal{D}$ not only as a regularizer, but as a mechanism for manifold expansion. By stochastically blocking latent paths during training, we force the model to distribute information across the entire rank spectrum, preventing the optimization from collapsing into a narrow subspace (rank collapse).
@@ 2.3 The Mergeability Trade-off
Therefore, the industrial standard is unmerged inference. In this paradigm, NoRA fits seamlessly. It requires no architectural changes to the serving infrastructure, only the execution of an additional activation kernel. We deliberately trade the theoretical mergeability—which is rarely used in multi-tenant scaling—for tangible gains in reasoning capacity.
@@ 2.4 Spectral Analysis Framework: Effective Rank
To quantify NoRA and LoRA’s “manifold expansion” capability, we adopt the effective rank (ER) metric [8]. Unlike the parametric rank (which is fixed), the effective rank measures the actual dimensionality of the information encoded in the activation space.
@@ 3.2 Datasets
  • MathInstruct (Secondary - Reasoning): A composite dataset of  100k mathematical problems (including GSM8K and MATH). We use this to validate NoRA’s domain adaptation capabilities, specifically its ability to model the complex logical dependencies required for mathematical derivation.
@@ 3.3 Baselines
We compare NoRA directly with the industry standard LoRA. To investigate rank saturation, we sweep across a logarithmic scale of ranks $r\in\{16,64,128,512\}$. This spectrum covers the transition from parameter-efficient constraints to high-capacity regimes.
@@ 4.1 The Capacity Scaling Law
We first investigate whether non-linearity allows the model to overcome the rank saturation observed in linear adapters. We train both LoRA and NoRA on the 300k-sample SlimOrca dataset in ranks $r\in\{16,64,128,512\}$.
Figure 2: The Capacity Scaling Law of SlimOrca. Although LoRA plateaus rapidly, NoRA continues to improve with rank. In particular, NoRA at rank 64 outperforms LoRA at rank 512, breaking the linear ceiling.
@@
Although the absolute PPL improvement from 3.90 (LoRA) to 3.83 (NoRA) demonstrates the benefit of non-linearity, the most critical finding is the crossover in efficiency. As shown in Figure 2, NoRA of rank 64 (PPL 3.89) outperforms LoRA of rank 512 (PPL 3.90). This indicates that NoRA achieves superior expressivity with $8\times$ fewer singular dimensions than the linear baseline, proving that the “linear ceiling” is a structural bottleneck rather than a parameter-capacity issue.
@@ 4.2 Generalization on Mathematical Reasoning
Figure 3: Scaling laws of MathInstruct. NoRA consistently outperforms LoRA across all ranks, with the gap widening at higher ranks ($r=512$). While LoRA continues to improve, it remains strictly dominated by NoRA’s non-linear adaptation curve.
@@
As illustrated in Figure 3, NoRA maintains a consistent advantage over LoRA throughout the rank spectrum. Mathematical derivation inherently requires capturing sharp, non-linear dependencies between logical steps; standard linear updates often struggle to model these rigid logical boundaries. By introducing non-linearity, NoRA effectively accommodates these complex mappings, achieving a lower perplexity of 1.97 at rank 512 compared to LoRA’s 2.07.
@@
Note that while NoRA dominates LoRA on MathInstruct, the performance gap is less pronounced than on SlimOrca. We attribute this to the difference in dataset diversity and manifold geometry. SlimOrca contains highly diverse, open-ended reasoning tasks (Chain-of-Thought, coding, common sense), which form a complex, irregular solution manifold that quickly saturates linear subspaces. In contrast, MathInstruct focuses on a more specialized domain where problem structures are more regular and patterned. In this highly structured manifold, linear approximations (LoRA) remain relatively effective for longer. Nevertheless, NoRA’s ability to squeeze out further gains confirms that even in structured domains, non-linearity provides a necessary edge for "last-mile" optimization.
@@ 4.3 Qualitative Analysis: Escaping the Linear Trap
Although perplexity and effective rank provide quantitative evidence of NoRA’s superiority, it is crucial to examine how this translates into actual reasoning capabilities. We conduct a qualitative case study using the MathInstruct dataset to analyze the generation behavior of both models.
@@
Table 1: Case study on iterative reasoning: Comparison of generation outputs for a logistic map problem. LoRA with a larger rank 512 suffers from state collapse, repeating the same value after the second step. In contrast, NoRA, even with a smaller rank 128, successfully models the non-linear recursive updates, generating a dynamic trajectory despite having $4\times$ fewer parameters.
@@
NoRA (Rank 128) – Dynamic Tracking
@@
  • NoRA (rank 128): In contrast, NoRA demonstrates robust dynamic tracking. It successfully recognizes that the value of $x_{n}$ must change at every step. Although NoRA utilizes a significantly smaller rank budget ($r=128$), its non-linear gating mechanism (SiLU) and the dropping mechanism allow the model to switch contexts and project the hidden states into a broader manifold.
@@
This qualitative difference aligns with our spectral analysis. The “heavy tail” in NoRA’s singular value spectrum reflects its ability to encode subtle state changes required for multi-step reasoning. LoRA’s rank collapse in the spectral domain manifests itself directly as repetitive generation in the text domain.
@@ 5 Mechanism & Design Analysis
In this section, we deconstruct NoRA to isolate the source of its performance. We move beyond predictive metrics (PPL) to analyze the structural properties of the learned weight manifolds using spectral analysis.
@@ 5.1 Spectral Signature of Non-linearity
To understand why NoRA utilizes parameters more efficiently than LoRA, we visualize the singular value spectrum of the learned adapter updates.
Figure 4: Spectral Signature of the SlimOrca dataset. LoRA exhibits “rank collapse,” where singular values drop precipitously, indicating under-utilization of the rank budget. NoRA maintains a “heavy tail,” activating a broader subspace.
@@ 5.1.1 Visualizing Expansion
As shown in Figure 4, standard LoRA in the SlimOrca dataset exhibits rank collapse: the singular values decay rapidly, effectively using only a fraction of the allocated rank. In contrast, NoRA demonstrates tail activation, maintaining significant energy across the singular value spectrum. This confirms that the non-linearity prevents the optimization path from being confined to a low-dimensional linear subspace.
@@
Figure 5: Effective Rank Analysis on SlimOrca. We compare the spectral utilization of LoRA and NoRA across increasing rank budgets ($r\in\{16,64,128,512\}$). As the rank budget increases to 512, LoRA’s effective rank saturates significantly, plateauing around $\sim$60. This confirms the “linear ceiling” hypothesis, where linear constraints prevent the model from utilizing the additional parameter budget. Conversely, NoRA demonstrates superior spectral efficiency. Its effective rank scales with the budget, reaching over 330 at $r=512$. This validates that our non-linear architecture successfully expands the representation manifold, avoiding the collapse observed in linear adapters.
@@ 5.1.2 Quantifying Manifold Expansion
We quantify the extent of subspace collapse using the effective rank (ER) metric [8]. As illustrated in Figure 5, a stark divergence emerges as the rank increases. At a target rank of $r=512$, LoRA suffers from severe spectral saturation, using only a fraction of its capacity with an effective rank of $\approx 60$. In contrast, NoRA successfully induces manifold expansion, maintaining a high effective rank that exceeds 330. This quantitative gap, where NoRA activates over $5\times$ more spectral dimensions than LoRA at high ranks, provides the geometric explanation for the efficiency crossover observed in Section 4.
@@
5.2 Deconstructing NoRA: An Ablation Study
@@
Table 2: Deconstructing NoRA (rank 128). We report test PPL (lower is better) on the SlimOrca dataset. The results confirm that the full NoRA architecture—combining weight-level granularity, SiLU gating, and structural dropout—yields the optimal performance.
@@
NoRA (Ours) Full Architecture 3.81
@@ 5.2.1 Granularity: The Weight-Level Advantage
Row (a) in Table 2 compares NoRA with a standard parallel adapter applied at the module output. The degradation in PPL (3.90) confirms that fine-grained intervention within the projection matrices ($W_{q},W_{v}$) is crucial. The Weight-level adaptation allows NoRA to manipulate the internal feature dynamics of the attention mechanism, whereas the module-level adaptation operates too coarsely.

### v2

# arXiv:2602.22911v2[cs.LG] 02 Mar 2026

@@ Abstract
Low-Rank Adaptation (LoRA) dominates parameter-efficient fine-tuning (PEFT). However, it faces a critical “linear ceiling” in complex reasoning tasks: simply increasing the rank yields diminishing returns due to intrinsic linear constraints. We introduce CeRA (Capacity-enhanced Rank Adaptation), a weight-level parallel adapter that injects SiLU gating and structural dropout to induce manifold expansion. On the SlimOrca benchmark, CeRA breaks this linear barrier: at rank 64 (PPL 3.89), it outperforms LoRA at rank 512 (PPL 3.90), demonstrating superior spectral efficiency. This advantage generalizes to mathematical reasoning, where CeRA achieves a perplexity of 1.97 on MathInstruct, significantly surpassing LoRA’s saturation point of 2.07. Mechanism analysis via Singular Value Decomposition (SVD) confirms that CeRA activates the dormant tail of the singular value spectrum, effectively preventing the rank collapse observed in linear methods.
@@ 1 Introduction
To break this ceiling, we introduce CeRA (Capacity-enhanced Rank Adaptation). CeRA marks a paradigm shift from linear subspace optimization to non-linear manifold deformation. By injecting SiLU gating and structural dropout directly into a fine-grained, parallel adapter architecture, CeRA unlocks the high-dimensional expressivity required for complex reasoning.
@@
  • Architecture: We propose CeRA, a fine-grained, weight-level parallel adapter that integrates non-linear gating to capture complex functional updates beyond linear approximations.
  • Empirical Scaling: We demonstrate that CeRA breaks the linear ceiling. On the large-scale SlimOrca benchmark [6], CeRA continues to scale, with CeRA at rank 64 significantly outperforming LoRA at Rank 512.
@@
  • Theoretical Mechanism: Through Singular Value Decomposition (SVD) analysis, we provide a spectral proof of manifold expansion. We show that CeRA activates the dormant tail of the singular value spectrum, preventing the rank collapse that constrains linear methods.
2 CeRA: Capacity-enhanced Rank Adaptation
We introduce CeRA, a methodology that reconciles the efficiency of parameter-efficient fine-tuning with the high-dimensional expressivity required for complex reasoning.
@@ 2.1 Preliminaries: The Linear Confinement
2.2 The CeRA Architecture
CeRA breaks this linear constraint. We retain the parallel bottleneck structure to preserve synchronization with the main branch but inject non-linearity to unlock high-rank capacity, as illustrated in Figure 1 (bottom-right). Formally, CeRA is defined as:
@@ 2.2 The CeRA Architecture
Figure 1: Architectural evolution of adaptation methods. Top Row (Module-Level): Traditional adapters operate on the output of the entire attention block. (Top-left) Serial Adapters create latency bottlenecks. (Top-right) Parallel Adapters improve efficiency but lack fine-grained control over internal projections. Bottom Row (Weight-Level): (Bottom-left) LoRA injects linear updates ($\Delta W=BA$) directly into $W_{q}$ and $W_{v}$. (Bottom-right) CeRA (Ours) operates at the same fine-grained level but introduces non-linear modeling via SiLU gating ($\sigma$) and structural Dropout ($\mathcal{D}$). This architecture combines the efficiency of weight-level injection with the high-capacity expressivity of non-linear manifolds.
@@
As shown in Figure 1, a critical distinction between CeRA and parallel adapters lies in the insertion point. Unlike traditional Parallel Adapters (top-right), which operate at the module level by processing the aggregate output of an attention block, CeRA operates at the weight level (bottom-right). By injecting updates directly into the internal query ($W_{q}$) and value ($W_{v}$) projections, CeRA fundamentally alters the internal feature dynamics of the attention mechanism rather than simply correcting its output.
@@
Linearity forces the adapter to process all input features uniformly. By introducing SiLU ($\sigma(x)=x\cdot\text{sigmoid}(x)$), CeRA gains a gating mechanism. This non-linearity allows the adapter to selectively suppress noise or amplify specific feature directions in the latent space, approximating complex decision boundaries that linear low-rank updates cannot represent.
@@
In standard PEFT, dropout is often omitted. In CeRA, we utilize dropout $\mathcal{D}$ not only as a regularizer, but as a mechanism for manifold expansion. By stochastically blocking latent paths during training, we force the model to distribute information across the entire rank spectrum, preventing the optimization from collapsing into a narrow subspace (rank collapse).
@@ 2.3 The Mergeability Trade-off
Therefore, the industrial standard is unmerged inference. In this paradigm, CeRA fits seamlessly. It requires no architectural changes to the serving infrastructure, only the execution of an additional activation kernel. We deliberately trade the theoretical mergeability—which is rarely used in multi-tenant scaling—for tangible gains in reasoning capacity.
@@ 2.4 Spectral Analysis Framework: Effective Rank
To quantify CeRA and LoRA’s “manifold expansion” capability, we adopt the effective rank (ER) metric [8]. Unlike the parametric rank (which is fixed), the effective rank measures the actual dimensionality of the information encoded in the activation space.
@@ 3.2 Datasets
  • MathInstruct (Secondary - Reasoning): A composite dataset of  100k mathematical problems (including GSM8K and MATH). We use this to validate CeRA’s domain adaptation capabilities, specifically its ability to model the complex logical dependencies required for mathematical derivation.
@@ 3.3 Baselines
We compare CeRA directly with the industry standard LoRA. To investigate rank saturation, we sweep across a logarithmic scale of ranks $r\in\{16,64,128,512\}$. This spectrum covers the transition from parameter-efficient constraints to high-capacity regimes.
@@ 4.1 The Capacity Scaling Law
We first investigate whether non-linearity allows the model to overcome the rank saturation observed in linear adapters. We train both LoRA and CeRA on the 300k-sample SlimOrca dataset in ranks $r\in\{16,64,128,512\}$.
Figure 2: The Capacity Scaling Law of SlimOrca. Although LoRA plateaus rapidly, CeRA continues to improve with rank. In particular, CeRA at rank 64 outperforms LoRA at rank 512, breaking the linear ceiling.
@@
Although the absolute PPL improvement from 3.90 (LoRA) to 3.83 (CeRA) demonstrates the benefit of non-linearity, the most critical finding is the crossover in efficiency. As shown in Figure 2, CeRA of rank 64 (PPL 3.89) outperforms LoRA of rank 512 (PPL 3.90). This indicates that CeRA achieves superior expressivity with $8\times$ fewer singular dimensions than the linear baseline, proving that the “linear ceiling” is a structural bottleneck rather than a parameter-capacity issue.
@@ 4.2 Generalization on Mathematical Reasoning
Figure 3: Scaling laws of MathInstruct. CeRA consistently outperforms LoRA across all ranks, with the gap widening at higher ranks ($r=512$). While LoRA continues to improve, it remains strictly dominated by CeRA’s non-linear adaptation curve.
@@
As illustrated in Figure 3, CeRA maintains a consistent advantage over LoRA throughout the rank spectrum. Mathematical derivation inherently requires capturing sharp, non-linear dependencies between logical steps; standard linear updates often struggle to model these rigid logical boundaries. By introducing non-linearity, CeRA effectively accommodates these complex mappings, achieving a lower perplexity of 1.97 at rank 512 compared to LoRA’s 2.07.
@@
Note that while CeRA dominates LoRA on MathInstruct, the performance gap is less pronounced than on SlimOrca. We attribute this to the difference in dataset diversity and manifold geometry. SlimOrca contains highly diverse, open-ended reasoning tasks (Chain-of-Thought, coding, common sense), which form a complex, irregular solution manifold that quickly saturates linear subspaces. In contrast, MathInstruct focuses on a more specialized domain where problem structures are more regular and patterned. In this highly structured manifold, linear approximations (LoRA) remain relatively effective for longer. Nevertheless, CeRA’s ability to squeeze out further gains confirms that even in structured domains, non-linearity provides a necessary edge for "last-mile" optimization.
@@ 4.3 Qualitative Analysis: Escaping the Linear Trap
Although perplexity and effective rank provide quantitative evidence of CeRA’s superiority, it is crucial to examine how this translates into actual reasoning capabilities. We conduct a qualitative case study using the MathInstruct dataset to analyze the generation behavior of both models.
@@
Table 1: Case study on iterative reasoning: Comparison of generation outputs for a logistic map problem. LoRA with a larger rank 512 suffers from state collapse, repeating the same value after the second step. In contrast, CeRA, even with a smaller rank 128, successfully models the non-linear recursive updates, generating a dynamic trajectory despite having $4\times$ fewer parameters.
@@
CeRA (Rank 128) – Dynamic Tracking
@@
  • CeRA (rank 128): In contrast, CeRA demonstrates robust dynamic tracking. It successfully recognizes that the value of $x_{n}$ must change at every step. Although CeRA utilizes a significantly smaller rank budget ($r=128$), its non-linear gating mechanism (SiLU) and the dropping mechanism allow the model to switch contexts and project the hidden states into a broader manifold.
@@
This qualitative difference aligns with our spectral analysis. The “heavy tail” in CeRA’s singular value spectrum reflects its ability to encode subtle state changes required for multi-step reasoning. LoRA’s rank collapse in the spectral domain manifests itself directly as repetitive generation in the text domain.
@@ 5 Mechanism & Design Analysis
In this section, we deconstruct CeRA to isolate the source of its performance. We move beyond predictive metrics (PPL) to analyze the structural properties of the learned weight manifolds using spectral analysis.
@@ 5.1 Spectral Signature of Non-linearity
To understand why CeRA utilizes parameters more efficiently than LoRA, we visualize the singular value spectrum of the learned adapter updates.
Figure 4: Spectral Signature of the SlimOrca dataset. LoRA exhibits “rank collapse,” where singular values drop precipitously, indicating under-utilization of the rank budget. CeRA maintains a “heavy tail,” activating a broader subspace.
@@ 5.1.1 Visualizing Expansion
As shown in Figure 4, standard LoRA in the SlimOrca dataset exhibits rank collapse: the singular values decay rapidly, effectively using only a fraction of the allocated rank. In contrast, CeRA demonstrates tail activation, maintaining significant energy across the singular value spectrum. This confirms that the non-linearity prevents the optimization path from being confined to a low-dimensional linear subspace.
@@
Figure 5: Effective Rank Analysis on SlimOrca. We compare the spectral utilization of LoRA and CeRA across increasing rank budgets ($r\in\{16,64,128,512\}$). As the rank budget increases to 512, LoRA’s effective rank saturates significantly, plateauing around $\sim$60. This confirms the “linear ceiling” hypothesis, where linear constraints prevent the model from utilizing the additional parameter budget. Conversely, CeRA demonstrates superior spectral efficiency. Its effective rank scales with the budget, reaching over 330 at $r=512$. This validates that our non-linear architecture successfully expands the representation manifold, avoiding the collapse observed in linear adapters.
@@ 5.1.2 Quantifying Manifold Expansion
We quantify the extent of subspace collapse using the effective rank (ER) metric [8]. As illustrated in Figure 5, a stark divergence emerges as the rank increases. At a target rank of $r=512$, LoRA suffers from severe spectral saturation, using only a fraction of its capacity with an effective rank of $\approx 60$. In contrast, CeRA successfully induces manifold expansion, maintaining a high effective rank that exceeds 330. This quantitative gap, where CeRA activates over $5\times$ more spectral dimensions than LoRA at high ranks, provides the geometric explanation for the efficiency crossover observed in Section 4.
@@
5.2 Deconstructing CeRA: An Ablation Study
@@
Table 2: Deconstructing CeRA (rank 128). We report test PPL (lower is better) on the SlimOrca dataset. The results confirm that the full CeRA architecture—combining weight-level granularity, SiLU gating, and structural dropout—yields the optimal performance.
@@
CeRA (Ours) Full Architecture 3.81
@@ 5.2.1 Granularity: The Weight-Level Advantage
Row (a) in Table 2 compares CeRA with a standard parallel adapter applied at the module output. The degradation in PPL (3.90) confirms that fine-grained intervention within the projection matrices ($W_{q},W_{v}$) is crucial. The Weight-level adaptation allows CeRA to manipulate the internal feature dynamics of the attention mechanism, whereas the module-level adaptation operates too coarsely.